In a problem of testing of a simple hypothesis against a simple alternative, if the probability of type 1 error is known to be 0.06 then the power of test will be0.94. Is it true?

1
Expert's answer
2021-12-27T04:26:57-0500

Solution:

It is false.



Given that "1-\alpha=0.06, 1-\beta=0.94".

Clearly, "\alpha, \beta" are unrelated. So, we cannot say that "\alpha=1-\beta".

Thus, false.